OpenStudy (anonymous):

Write the quadratic equation whose roots are -1 and 4, and whose leading coefficient is 1. (Use the letter to represent the variable.)

OpenStudy (anonymous):

x

jimthompson5910 (jim_thompson5910):

x = -1 or x = 4 x + 1 = 0 or x - 4 = 0 (x+1)(x-4) = 0 x^2-3x-4=0

OpenStudy (anonymous):

thank you that was correct so if the numbers were switched out and it was roots -1 and 4 with leading coefficent of 1

OpenStudy (anonymous):

would it be written as x=-3 or x=4 x-3=0 or x+6=0 (x-3)(x+6)=0

jimthompson5910 (jim_thompson5910):

wait, is this a different problem?

OpenStudy (anonymous):

yesss

jimthompson5910 (jim_thompson5910):

what are the roots again?

OpenStudy (anonymous):

x=-3 or x=6 coefficent 1

jimthompson5910 (jim_thompson5910):

ok thx

jimthompson5910 (jim_thompson5910):

x=-3 or x=6 x+3=0 or x-6=0 (x+3)(x-6)=0 x^2-3x-18=0

OpenStudy (anonymous):

so the first one is always postive 2nd always negative

jimthompson5910 (jim_thompson5910):

well because the roots have different signs, the factors have different signs

OpenStudy (anonymous):

so one with postive roots 4,5 with cofficent of 1

jimthompson5910 (jim_thompson5910):

x=4 or x=5 x-4=0 or x-5=0 (x-4)(x-5)=0 x^2-9x+20=0

OpenStudy (anonymous):

dang your good

jimthompson5910 (jim_thompson5910):

does that answer your question?

OpenStudy (anonymous):

thank you really appreciate your help

jimthompson5910 (jim_thompson5910):

lol thx, glad to be of help